A circular antenna of area 3m² is installed at a particular place. The plane of the area of antenna is inclined at 47⁰ with the direction of Earth's magnetic field. If the magnitude of Earth's field at that place is 40773.9 nT, find the magnetic flux linked with the antenna.

1
Expert's answer
2022-03-16T18:28:42-0400

The magnetic flux linked with the antenna

"\Phi=BA\cos\theta\\
={\rm 40773.9*10^{-9}\: T*3.0\: m^2*\cos 43^\circ}\\
=\rm 8.95*10^{-5}\; Wb"